OpenStudy (anonymous):

-10X-16X+18X=60-26X+8X

OpenStudy (anonymous):

Use the method I taught you.

OpenStudy (anonymous):

Does this involve about combining like terms?

OpenStudy (anonymous):

combine all of your like terms and then solve for x. The answer should come out to be x = 6

OpenStudy (anonymous):

i DONT WANT TO SEEM SO DUMB BUT WITH MATH YES I AM i HAVE SEVERE ISSUES WITH PLEASE EXUSE ME FOR ASKING FOR SO MUCH HELP BUT I NEED ALL THE HELP i CAN GET SORRY

OpenStudy (anonymous):

-10X-16X+18X=60-26X+8X combine like terms (everything with an x) on the left and right side -8x = 60 -18x add 18x to the left side 10x = 60 divide by 10 to isolate x x = 6

OpenStudy (anonymous):

THANK YOU FOR YOU STEP BY STEP ANSWER I WILL NOT BOTHER ANY MORE YOU HAVE HELPED ME SO MUCH YOU JUST DONT KNOW I APPRECIATE ,THE WAY YOU RESPONDED TO MY PROBLEM YOUR A VERY HELPFUL/KIND PERSON TO BE SO PATIENET AND NOT JUDGE MY STUPIETY

OpenStudy (anonymous):

No problem! I'm a college student currently taking cal 3, and I enjoy helping other people with their math so I can avoid doing my own. Its a very strange form of procrastination. Hope I was able to help!

OpenStudy (anonymous):

you do not even know you were so helpful I appreciate it Have a fantastic weekend.Alexias

OpenStudy (anonymous):

\[-10x-16x+18x \rightarrow 60-26x+8x , -26x+18x \rightarrow 60-26x+8x , -26x+26x+18x-8x \rightarrow60 , 10x \rightarrow 60 , x \rightarrow 60/10 , ans: x \rightarrow 6\]